You are given the following information about a country:
Billions ($)
Receipts of factor income from the rest of the world 25
Indirect taxes 60
Corporate profits 80
Net interest 15
Depreciation 150
Compensation of employees 850
Net business transfer payments 20
Amount of national income not going to households 100
Proprietor’s income 220
Personal income taxes 200
Rental income 35
Subsidies 30
Payments of factor income to the rest of the world 45
Surplus of Governmental Enterprises 50
Using the information provided on the table above to compute the following (Show your work
and formulas).
a. National Income
b. Gross National Product
c. Gross Domestic Product
d. Disposable Personal Income
Please don't forget the formulas. thank you
1
Expert's answer
2020-05-08T14:21:03-0400

a.National Income=Gross National Product-Depreciation-Indirect taxes=1410-150-60=1200

b.Gross National Product=Gross Domestic Product-(Receipts of factor income from the rest of the world -Payments of factor)+ income to the rest of the world)=1390-(25-45)=1410

c.Gross Domestic Product=(Receipts of factor income from the rest of the world -Payments of factor)+ income to the rest of the world)+Indirect taxes+Corporate profits+Net interest+ Depreciation+Compensation of employees+Net business transfer payments+Amount of national income not going to households+Proprietor’s income-Personal income taxes+Rental income-Subsidies+Surplus of Governmental Enterprises=(25-45)+60+80+15+150+850+20+100+220-200+35-30+50=1390

d. Disposable Personal Income=National Income-Amount of national income not going to households-Personal income taxes=1200-100-200=900